Find the angle between the vectors √2i + 2j + 2k and i + √2j + √2k.


1
Expert's answer
2022-01-30T15:19:27-0500

Solution:


Let "\overrightharpoon{a}=\sqrt{2}i+2j+2k" and "\overrightharpoon{b}=i+\sqrt{2}j+\sqrt{2}k."


Now angle between "\overrightharpoon{a}" and "\overrightharpoon{b}" be:



"cos(\phi)=(\frac{\overrightharpoon{a}\overrightharpoon{b}}{|\overrightharpoon{a}||\overrightharpoon{b}|})=(\frac{5\sqrt{2}}{(\sqrt{10}(\sqrt{5}))})=1"

(in case of complex numbers, we need to take the real part of the dot product).


"\phi = acos (1) = 0 = 0^o"